OpenStudy (anonymous):

5r-4s-3t=3 t=s+r r=3s+1

OpenStudy (anonymous):

s this three sets of equations? solving for all three variables? 5r-4s-3t=3 t=s+r r=3s+1 They are pretty crammed together, I need to know what I'm starting with and where i'm going/

OpenStudy (anonymous):

{5r-4s-3t=3 {t=s+r {r=3s+1 HELP?!

OpenStudy (anonymous):

t=s+r s=(r-1)/3 r=3s+1 all term in s and r lots of math later s=9/7 r=15/7